A solid ball is made of an insulating material that has charge uniformly distributed throughout its volume, not just on its surface. The ball has a radius of 26 cm and it has a total charge of 24 HC Find the electric field strength at the following distances from the balls center (a) 21 cm (This is inside the ball) kN/C (b) 26 cm (Right at the surface) kN/C )52 cm (Outside the bal) kN/C
Answer #1

2 r3 1 Qr (field inside a uniformly charged sphere)

1 2 4πε0,2 (field outside a uniformly charged sphere)

a) E= 9*10^9 * 24*10^-6* 0.21/0.26^3
E = 2580.79 kN/C

b) E= 9*10^9 * 24*10^-6/0.26^2
E = 3195.266 kN/C

c) E = 9*10^9 * 24*10^-6/0.52^2
E = 798.81 kN/C